OpenStudy (anonymous):

Two square roots for 121(cos315° + i sin315°) are ?

OpenStudy (anonymous):

Please help me :|

OpenStudy (anonymous):

121^(1/2) to get 11 faster :)

OpenStudy (anonymous):

but how to keep going from there?

OpenStudy (anonymous):

i mean we knows that : z = 11 cis (157.5degrees + 180degrees k) how to find w0 and w1 ?

OpenStudy (anonymous):

hmmm in class i learned a totally different way compare to what you did here ^^

OpenStudy (anonymous):

let me write it out and you see it, because i don't get it :(

OpenStudy (anonymous):

so with this example : The 3 cube roots of z = 8(cos60°+ i sin60°)

OpenStudy (anonymous):

are you looking for the square root of \[121\left(\cos(315)+i\sin(315)\right)\]?

OpenStudy (anonymous):

i don't think so no

OpenStudy (anonymous):

that looks like the question you asked yes?

OpenStudy (anonymous):

or was the original question something different?

OpenStudy (anonymous):

step 1 : wk = 8^1/3 cis (60°/3 + 360° k) according to the formula

OpenStudy (anonymous):

step 2 : =2 cis(20°+120°k)

OpenStudy (anonymous):

what was the original question for this ??

OpenStudy (anonymous):

where k = 0,1,2. and they start to find w0 =, w1=,w2= .... and i don't get how they do that

OpenStudy (anonymous):

Two square roots for 121(cos315° + i sin315°) are ? <---- this was original

OpenStudy (anonymous):

I know that I am right till here--> z=121(1/sqrt(2) - i(1/sqrt(2)) )

OpenStudy (anonymous):

ok so we can do that easy first take the square root of 121, get 11 then divide the angle by 2 get \[11\left(\cos(157.5)+i\sin(157.5)\right)\] for one answer. the other answer is half way around the unit circle

OpenStudy (anonymous):

yes yes, i got the samething yay! :) the other answer is 180

OpenStudy (anonymous):

couldn't you just put it together for a simpler form like : 11 cis (157.5 + 180k) ?

OpenStudy (anonymous):

so start again with \[121\left(\cos(315)+i\sin(315)\right)\] add 360 to the angle get \[121\left(\cos(675)+i\sin(675)\right)\]take half of that angle and get your other solution \[11\left(\cos(337.5)+i\sin(337.5)\right)\]

OpenStudy (anonymous):

wait what? =.=

OpenStudy (anonymous):

sure you can do what you wrote, but there are only two

OpenStudy (anonymous):

i thought it was 180

OpenStudy (anonymous):

it is not 180 it is what you wrote, \(157.5+180\)

OpenStudy (anonymous):

ohh so you added them together

OpenStudy (anonymous):

this is what you wrote 11 cis (157.5 + 180k) and you only need k = 0 which gives 157.5 and k = 1 which gives 337.5

OpenStudy (anonymous):

after that you are just going around the circle again so all the numbers repeat. there are only two square roots

OpenStudy (anonymous):

wait wait, i think the answer need to be in trigonometric form

OpenStudy (anonymous):

i wrote the answer in trig form above. there are two square roots. \[11\left(\cos(157.5)+i\sin(157.5)\right)\] and \[11\left(\cos(337.5)+i\sin(337.5)\right)\]

OpenStudy (anonymous):

those are the answers ?

OpenStudy (anonymous):

@satellite73 , he is looking for the values

OpenStudy (anonymous):

yes those are the two answers

OpenStudy (anonymous):

I think he is looking for this method --> http://answers.yahoo.com/question/index?qid=20081102074555AAAJbwx

OpenStudy (anonymous):

oh lol, what about if there were 3 answers? what do you do to find the third one ?

OpenStudy (anonymous):

@shivam_bhalla i doubt that. there is no way in the world he is going to be able to evaluate those, unless you are going to use some half angle formulas. it is doable, but he said "trig form" so those are the answers

OpenStudy (anonymous):

@Ala123 there will be three answers if you are looking for cube roots

OpenStudy (anonymous):

@shivam_bhalla : satellite73 method is just like in my book ^^

OpenStudy (anonymous):

4 answers for 4th roots, etc

OpenStudy (anonymous):

Great. Good work :)

OpenStudy (anonymous):

cool, i get it now thanks guys !

OpenStudy (anonymous):

now let me try cube root and i'll be back asking for help again =.=